OpenStudy (anonymous):

In Addition to paying $185 per unit ,a community college student pays a registration fees of $200 per semester .The equation that gives the cost (y)to take x units is y=185x+200 (A)Find the y-intercept of this equation .In the context of this problem,is this cost possible ?Explain. (B)In your own words,explain why this equation has no x-intercept in the context of this problem. (C)A full load for a student is 12units. How much will a full time student pay per semester

OpenStudy (anonymous):

The y-intercept is the "b" in the equation y=mx+b. y=185x+200 The y-intercept is the fixed cost of $200. The registration fee includes $200 whether you take one credit or ten.

OpenStudy (anonymous):

B. I sort of answered this already...you have a $200 registration fee no matter what. It's the fixed cost. C. A full time student taking 12 units of credit--x is the variable for cost (it's dependent on the number of credits being taken). a full time student-----y=185(12)+200=2220 + 200 = $2420. Hope this helps